The base of a triangular sail is 16 m shorter than its height. The area is 130m squared. Find the base and height of the sail.
Answer by Cromlix(4381)

Hi there,
Area of triangle = 1/2(base x height)
Make height = x
Base = x - 16
Area = 1/2(base x height)
130 = 1/2((x - 16) * x) (* times)
130 = 1/2(x^2 - 16x)
130 = x^2/2 - 8x
Put into ax^2 + bx + c = 0 form
x^2/2 - 8x - 130 = 0
Multiply through by 2
x^2 - 16x - 260 = 0
Factorise:
(x + 10)(x - 26)= 0
x + 10 = 0
x = -10 (disregard as -ve)
x - 26 = 0
x = 26
Height of sail = 26m
Base of sail = 10m
Hope this helps :-)